To determine the number of deer in a game preserve, a conservationist catches 375 deer, tags them and lety them loose. Later, 732 deer are caught; 183 of them are tsgged. How many deer are in the preserve?
How many deer are in the preserve?
____deer

Answer by stanbon(57967)

To determine the number of deer in a game preserve, a conservationist catches 375 deer, tags them and lets them loose. Later, 732 deer are caught; 183 of them are tsgged. How many deer are in the preserve?
How many deer are in the preserve? ____deer
-----------------------------------------------------
pop/tagged = captured/tagged when captured
---
x/375 = 732/183
x = 375(732/183) = 1500 (population estimate)
----
Cheers,
Stan H.